OpenStudy (anonymous):

order these units from smallest to largest:cm,um,km,mm,m,nm,dm,pm. then give each a measurement in terms of meters.

OpenStudy (anonymous):

pm, nm, um, mm, cm, dm, m, km

OpenStudy (anonymous):

Welcome to OS :3

OpenStudy (anonymous):

megameter= 10^6 meters= 1,000,000 meters kilometer= 10^3 meters=1000 meters hectometer= 10^2 meters=100 meters decameter= 10^1 meters=10 meters decimeter= 10^−1 meters=0.1 meters centimeter= 10^−2 meters=0.01 meters millimetre 10^−3 metres micrometers= 10^−6 meters=0.000 001 meters nanometers= 10^−9 meters=0.000 000 001 meters picometers= 10^−12 meters= 0.000 000 000 001 meters